What amount of sodium dodecyl sulfate (SDS) do you need to weight out on a balance to prepare 550 mL of a 30 % (w/v) solution?
Answer #1

Answer :

30% (w/v) Solution means 30 g of solute is present in 100 ml of solution .

Here solute is sodium dodecyl sulfate (SDS) .

Total volume of solution = 550 ml (given)

Since 100 ml of solution contain = 30 g of solute

1 ml of solution contain = 30 / 100 = 0.3 g of solute

So, 550 ml of solution contain = 0.3 * 550 = 165 g of solute

Amount of sodium dodecyl sulfate required is 165 g .